OpenStudy (anonymous):

We toss 3 coins 100,000 times. How many times do we expect to see "Exactly 1 Head" appearing? A. 12,500 B. 37,500 C. 3,750 D. 50,000

OpenStudy (anonymous):

C

OpenStudy (anonymous):

150,000?

OpenStudy (anonymous):

3coins(50,000times per coin)?

OpenStudy (anonymous):

hold on i've calculated how many times 3 heads would appear on each toss. is that whats required???

OpenStudy (anonymous):

You are flipping 3 coins at the same time, and you are doing this 100000 times. So you could get things like: HHH TTH THT etc. The question is how many times do you expect exactly one heads to show up. So you want TTH THT HTT

OpenStudy (turingtest):

possible configurations with one head upward:3 total possible configurations of the three coins=2^3=8 so each toss has a 3/8 chance of puling up one heads right guys? (I'm not the best with probability0

OpenStudy (anonymous):

There are 8 possible ways for the coins to turn up (2 for each coin, so 2^3 = 8), and three of those possibilities have exactly one head. So the probability is 3/8. 3/8 of 100000 is 37500. I think this is the correct way to interpret the problem.

OpenStudy (dumbcow):

joemath is correct

OpenStudy (turingtest):

Yes, I believe Joemath's answer is the correct one.

OpenStudy (anonymous):

yes - i agree